# Non-Adaptive Multi-Stage Algorithm and Bounds for Group Testing with Prior Statistics

### v4

@@ III-A1 Testing Matrix and Pooling
The proposed multi-stage recovery algorithm is intended to work with any non-adaptive testing matrix, e.g., as given in [46]. To simplify the technical aspects and focus on the key methods, the testing matrix is generated randomly under a fixed optimal approximation with Bernoulli distribution of $p=\ln(2)/K$ [47], using classical GT methods. The pooling and its outcome are given by the process elaborated in Section II and illustrated in Fig. 1.
Figure 1: For an unknown population $\mathbf{U}\in\left\{0,1\right\}^{9}$ with $K=2$, a random testing matrix is sampled and the test result $Y$ is calculated.
Algorithm 1 Multi-Stage Recovery Algorithm
Input: $\mathbf{X},Y,\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N},K,L,\tau,\gamma$
@@
2:$\mathcal{P}^{(DND)}\leftarrow\mathbf{DND}(\mathbf{X},Y)$
3:$\mathcal{P}^{(DD)}\leftarrow\mathbf{DD}\left(\mathbf{X},Y,\mathcal{P}^{(DND)}\right)$
4:$\mathcal{P}^{(S_{1,2})}\leftarrow\mathcal{N}\setminus\left(\mathcal{P}^{(DND)}\cup\mathcal{P}^{(DD)}\right)$
@@
6:$\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N}\leftarrow\mathbf{{updatePriors}}\left(\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N},\mathcal{P}^{(S_{1,2})},\mathcal{P}^{(DD)}\right)$
7:$\mathbf{Z}\leftarrow\mathbf{LVA}\left(L,\tau,\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N}\right)$
8:$\mathbf{C}\leftarrow\{\}$
@@
10:    if $K\leq\sum_{i}\left(\mathbf{Z}_{l,i}\right)\leq\gamma K$ then
@@
12:         $\mathbf{C}\leftarrow\mathbf{C}\cup\mathbf{getAllCombinations}\left(\mathbf{V}^{(l)},K\right)$
@@
15:$\hat{\mathcal{K}}\leftarrow\mathbf{MAP}\left(\mathbf{X},Y,\mathbf{C},\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N}\right)$
@@ III-A2 Recovery Process
The suggested recovery algorithm operates in two main stages. In the first stage (Stage 1), to reduce the space of search (i.e., the possible defective items), the algorithm efficiently identifies non-defective and definitely defective items without considering the prior correlated information. In the first step of this stage, we use the Definitely Not Defective (DND) algorithm [48, 49, 50, 19] (line 2, Fig. 2.(a)). DND compares the columns of the testing matrix, $\mathbf{X}$, with the outcome vector, $Y^{T}$. If $Y(i)=0$ for some $i\in\{1,\ldots,T\}$, the algorithm eliminates all items participating in the $i$-th test from being defective, and outputs them as the set $\mathcal{P}^{(DND)}\subset\mathcal{N}$.
@@
In the second step of Stage 1, we use the Definitely Defectives (DD) algorithm [19] (line 3, Fig. 2.(b)), which goes over the testing matrix and the test result, looking for positive pool tests that include only one possibly defective item. DD denotes those items as definitely defective items and outputs them as the set $\mathcal{P}^{(DD)}$.
@@
Let $\mathcal{P}^{(S_{1,1})}=\mathcal{N}\setminus\mathcal{P}^{(DND)}$ and $\mathcal{P}^{(S_{1,2})}=\mathcal{N}\setminus\left(\mathcal{P}^{(DND)}\cup\mathcal{P}^{(DD)}\right)$ denote the set of items that their status is still unclear after the first step and the second step, respectively. $\mathcal{P}^{(S_{1,2})}$ holds a new space search, and $\mathcal{P}^{(DD)}$ holds the already known defectives. This knowledge is acquired without utilizing any prior data, which we reserve for the second stage.
@@
Figure 2: First stage of MSGT. (a) The first step of Stage 1, the DND algorithm, reveals 5 DND items in $\mathbf{U}$, forming $\mathcal{P}^{(DND)}$. Since items participating in negative tests must be non-defective, we mark all the participants in the two negative test results as non-defective. (b) The second step of Stage 1, the DD algorithm, outputs $\mathcal{P}^{(DD)}$ that includes a single DD item, based on the first test result, as it is the only possibly defective item participating in this test. The two other positive test results do not contribute to our knowledge here because there is more than one possibly defective item participating in them.
@@
In the first step of Stage 2, we translate the data we obtained in DND and DD, $\mathcal{P}^{(S_{1,2})}$ and $\mathcal{P}^{(DD)}$, into the state space in terms of transition probabilities, $\{\mathbf{\Phi}_{i}\}_{i=1}^{N}$, and initial probabilities, $\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N}$, so we can employ all the gathered information in the next steps (line 6, Fig. 3.(a)). In the state space, the population sequence, $\mathbf{U}$, is parallel to time steps considered traditionally in Viterbi Algorithm (VA) for communications problems [32], and there are two possible states per item, the first indicates “non-defective” and the second indicates “defective”.
@@
In the second step, the suggested LVA for GT goes over the sequence of items and outputs $\mathbf{Z}\in\{0,1\}^{L\times N}$, a list of the $L$ most likely trajectories in the state space (line 7, Fig. 3.(b)) according to MAP decision based on the given prior information. Each trajectory is a sequence of $N$ states, representing items classified as either defective or non-defective. Thus, $\mathbf{Z}$ provides $L$ estimations of $\mathbf{U}$. In practice, the $L$ estimations may include any number of defective items and require further processing.
@@
In the third step, we extract candidates for the defective set out of the $L$ estimated sequences $\mathbf{Z}$ (lines 9 to 14). For some $l\in\{1,\ldots,L\}$, let $\mathbf{V}^{(l)}$ denote the set of items estimated as defective in $\mathbf{Z}_{l}$, the $l$-th row of $\mathbf{Z}$ (line 11). We ignore sequences that contain less than $K$ defective items or more than $\gamma K$ defective items, for some $\gamma\geq 1$, and consider only $\mathbf{Z}_{l}$ in which $K\leq\left|\mathbf{V}^{(l)}\right|<\gamma K$ as valid sequences. For each one of the valid sequences, we refer to all the combinations of size $K$ in $\mathbf{V}^{(l)}$ as candidates for $\hat{\mathcal{K}}$, and add them to the candidates list $\mathbf{C}$ (line 12).
@@
At this point, we have in $\mathbf{C}$ a list of candidates to be our final estimation $\hat{\mathcal{K}}$, and we can calculate the probability of each one of them using $\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N}$. Then, in the fourth step, the estimated defective set, $\hat{\mathcal{K}}$, is finally chosen using MAP estimator out of the $\mathbf{C}$ (line 15), i.e., $\hat{\mathcal{K}}=\arg\max_{c\in\mathbf{C}}{P\left(Y\lvert\mathbf{X},c\right)P\left(c\right)}$.
@@
If there are no valid sequences in $\mathbf{Z}$, we consider trajectories with fewer than $K$ detections for partial recovery. We select the trajectory with the most detections and randomly complete it to form a set of size $K$ for our final estimation $\hat{\mathcal{K}}$.
@@
Figure 3: Stage 2 of MSGT. (a) All the possible transitions in the state space that we consider in the LVA step, following the insights obtained in Stage 1. These transitions aggregate to a total of 6 trajectories. (b) The two most likely trajectories returned by LVA (assuming $L=2$). Given $K=2$, the black trajectory corresponds to a valid population vector $\mathbf{U}$ with 2 defective items, while the gray trajectory indicates an invalid population with 3 defective items instead. Consequently, in the subsequent step, MSGT will extract two optional defective sets: $\left\{U_{6},U_{8}\right\}$ and $\left\{U_{7},U_{8}\right\}$, and will finally choose the most likely one using MAP estimator. (c) Comparison of Stage 2 to ML. With $T=5$, we use the first 5 rows of the testing matrix, ignoring the last test result. This leaves 3 possibly defective items, forming two potentially defective sets of size $K=2$. Using ML, one set is chosen randomly, leading to an error probability of 0.5. With $T=6$, based on the third and sixth test results, there is only one set of size $K=2$ that matches the outcome $Y$, resulting in successful decoding with the ML decoder. As shown above, MSGT’s Stage 2 can successfully decode $\mathbf{U}$ with just $T=5$, as using the LVA step it narrows down to only 2 possible trajectories, and then the final estimation is selected based on the given prior information and the insights gained in Stage 1.
@@ III-B Analytical Results
Here we provide analytical results related to the proposed MSGT.
Let $P_{e,a}^{(DND)}$ and $P_{e,u}^{(DND)}$ denote the error probability of DND in the average case and its deviation from the average (“worst-case”) that still allows the success of MSGT on average in feasible computational complexity, respectively. The two following theorems give the upper bound on the expected number of possible defective items and a “worst-case” upper bound after the first step of Stage 1.
Theorem 1 (Cohen et al. [4]).
Consider a group test with a Bernoulli testing matrix with $p=\ln{2}/K$, and $T$ tests as $K\rightarrow\infty$. Let $P_{e,a}^{(DND)}\triangleq N^{-\alpha\left(1-\ln{2}/K\right)/2}$ for $\alpha\triangleq T/K\log_{2}{N}$. The expected number of possibly defective items is bounded by
$\mathbb{E}\left[\left|\mathcal{P}^{(S_{1,1})}\right|\right]\leq K+\left(N-K\right)P_{e,a}^{(DND)}.$
@@
According to [51, Lemma 2]:
$\mathbb{E}\left[\left|\mathcal{P}^{(S_{1,1})}\right|\right]=K+\left(N-K\right)P_{e}^{(DND)},$ (1)
where $P_{e}^{(DND)}=\left(1-p\left(1-p\right)^{K}\right)^{T}$ is the probability of non-defective item to be hidden by another defective item during DND.
By substituting $T=\alpha K\log_{2}{N}$, $p=\ln{2}/K$ we bound the error probability as follows:
$P_{e}^{\left(DND\right)}=\left(1-\frac{\ln{2}}{K}\left(1-\frac{\ln{2}}{K}\right)^{K}\right)^{\alpha K\log_{2}{N}}$
$=\left(1-\frac{\ln{2}}{K}\left(1-\frac{\ln{2}}{K}\right)^{\left(K-1\right)}\left(1-\frac{\ln{2}}{K}\right)\right)^{\alpha K\log_{2}{N}}$
$\leq\left(1-\frac{\ln{2}}{K}e^{-\ln{2}}\left(1-\frac{\ln{2}}{K}\right)\right)^{\alpha K\log_{2}{N}}$
$=e^{-\frac{\alpha}{2}\ln{2}\left(1-\frac{\ln{2}}{K}\right)\log_{2}{N}}$
$=N^{-\frac{\alpha}{2}\left(1-\frac{\ln{2}}{K}\right)}$
Finally, we denote this bound on the average error probability in DND by $P_{e,a}^{(DND)}\triangleq N^{-\frac{\alpha}{2}\left(1-\ln{2}/K\right)}$. Hence, substituting $P_{e,a}^{(DND)}$ in (1), we have:
$\mathbb{E}\left[\left|\mathcal{P}^{(DND)}\right|\right]\leq K+\left(N-K\right)P_{e,a}^{(DND)}.$
This completed the theorem proof. ∎
@@
Consider a group test with a Bernoulli testing matrix and DND decoder. As $K\rightarrow\infty$, the worst-case error probability of DND is bounded by
$P_{e,u}^{(DND)}$ $\leq P_{e,a}^{(DND)}+\left(1+\frac{\left(\gamma K-\left(K+\left(N-K\right)P_{e,a}^{(DND)}\right)\right)^{2}}{\left(N-K\right)P_{e,a}^{(DND)}\left(1-P_{e,a}^{(DND)}\right)}\right)^{-1},$
for any $\gamma\geq 1$.

### v5

@@ III-A1 Testing Matrix and Pooling
The proposed two-stage recovery algorithm is intended to work with any non-adaptive testing matrix, e.g., as given in [37]. To simplify the technical aspects and focus on the key methods, the testing matrix is generated randomly under design sparsity bounds $(K_{\min},K_{\max})$ (as defined in Section II) with Bernoulli distribution of $p=\ln(2)/K_{\max}$ [2], using classical GT methods. The pooling and its outcome are given by the process elaborated in Section II and illustrated in Fig. 2. We emphasize that the testing matrix is intentionally chosen to be a standard Bernoulli design, so that the impact of exploiting correlated prior information is isolated to the decoding stage [^3] An interesting direction for future work is the design of non-adaptive testing matrices that explicitly exploit non-uniform or correlated priors, for example via item-dependent testing probabilities..
Figure 2: For an unknown population $\mathbf{U}\in\left\{0,1\right\}^{9}$ a realized defective set of size $K=2$ and design interval $K_{\min},K_{\max}]=[2,2]$, a random testing matrix is sampled and the test result $Y$ is calculated.
Algorithm 1 Two-Stage Recovery Algorithm
Input: $\mathbf{X},Y,{\color[rgb]{0,0,0}\mathbf{\pi},\{\mathbf{\Phi}_{i}\}_{i=\tau+1}^{N}},{\color[rgb]{0,0,0}K_{\min},K_{\max}},L,\tau,\gamma$
@@
[…]
@@
[…]
@@
10:  if ${\color[rgb]{0,0,0}K_{\min}\leq}\sum_{i}\left(\mathbf{Z}_{l,i}\right)\leq{\color[rgb]{0,0,0}\lfloor}\gamma{\color[rgb]{0,0,0}K_{\max}}{\color[rgb]{0,0,0}\rfloor}$ then
@@
12:    $\mathbf{C}\leftarrow\mathbf{C}\cup\mathbf{getAllCombinations}\left(\mathbf{V}^{(l)},{\color[rgb]{0,0,0}K_{\min}},{\color[rgb]{0,0,0}K_{\max}}\right)$
@@
[…]
@@ III-A2 Recovery Process
The suggested recovery algorithm operates in two main stages. In the first stage (Stage 1), to reduce the space of search (i.e., the possible defective items), the algorithm efficiently identifies non-defective and definitely defective items without considering the prior correlated information. In the first step of this stage, we use the Definitely Not Defective (DND) algorithm [38, 16, 15, 3] (line 2, Fig. 3.(a), and Algorithm 2). DND compares the columns of the testing matrix, $\mathbf{X}$, with the outcome vector, $Y^{T}$. If $Y(i)=0$ for some $i\in\{1,\ldots,T\}$, the algorithm eliminates all items participating in the $i$-th test from being defective, and outputs them as the set $\mathcal{P}^{(DND)}\subset\mathcal{N}$.
@@
In the second step of Stage 1, we use the Definitely Defectives (DD) algorithm [3] (line 3, Fig. 3.(b), and Algorithm 3), which goes over the testing matrix and the test result, looking for positive pooled tests that include only one possibly defective item. DD denotes those items as definitely defective items and outputs them as the set $\mathcal{P}^{(DD)}$.
@@
Let $\mathcal{P}^{(S_{1,1})}=\mathcal{N}\setminus\mathcal{P}^{(DND)}$ and $\mathcal{P}^{(S_{1,2})}=\mathcal{N}\setminus\left(\mathcal{P}^{(DND)}\cup\mathcal{P}^{(DD)}\right)$ denote the set of items that their status is still unclear after the first step and the second step, respectively, where we use the notation $S_{i,j}$ with the first index $i$ indicating the stage and the second index $j$ indicating the step within that stage. $\mathcal{P}^{(S_{1,2})}$ holds a new space search, and $\mathcal{P}^{(DD)}$ holds the already known defectives. This knowledge is acquired without utilizing any prior data, which we reserve for the second stage.
@@
Figure 3: First stage of 2SDGT. (a) The first step of Stage 1, the DND algorithm, reveals 5 DND items in $\mathbf{U}$, forming $\mathcal{P}^{(DND)}$. Since items participating in negative tests must be non-defective, we mark all the participants in the two negative test results as non-defective. (b) The second step of Stage 1, the DD algorithm, outputs $\mathcal{P}^{(DD)}$ that includes a single $D$ item, based on the first test result, as it is the only possibly defective item participating in this test. The two other positive test results do not contribute to our knowledge here because there is more than one possibly defective item participating in them.
@@
In the first step of Stage 2, we translate the data we obtained in DND and DD, $\mathcal{P}^{(S_{1,2})}$ and $\mathcal{P}^{(DD)}$, into the state space in terms of transition probabilities, $\color[rgb]{0,0,0}\mathbf{\pi}$}, and initial probabilities, $\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N}$, so we can employ all the gathered information in the next steps (line 6, Fig. 4.(a), and Algorithm 4). In the state space, the population sequence, $\mathbf{U}$, considered herein, is replaced with time steps considered traditionally in VA for communications problems [56], in which we have in GT two possible states per item, the first indicates “non-defective” and the second indicates “defective”.
Remark 1.
Under the Markovian prior model described in Section II, the population vector $\mathbf{U}=(U_{1},\ldots,U_{N})$ admits a natural representation as a path in a trellis, where each state corresponds to the defectivity status of an item (or a $\tau$-length state for $\tau$-order memory), and the transition probabilities are given by the prior parameters $(\pi,\{\Phi_{i}\})$. In this representation, exact MAP decoding corresponds to finding the most likely path through the trellis given the observed test outcomes. However, exact MAP decoding over this trellis requires evaluating an exponential number of paths and is therefore computationally infeasible for large $N$. The LVA provides an efficient approximation by returning a list of the $L$ most likely trellis trajectories, rather than committing to a single path. This is particularly well suited to the proposed framework, since after the Stage 1 reduction (DND and DD), multiple high-probability trajectories may remain consistent with the observations. By producing a small set of candidate trajectories, LVA enables systematic exploitation of the correlated prior structure while maintaining feasible computational complexity.
@@
In the second step, the suggested LVA for GT goes over the sequence of items and outputs $\mathbf{Z}\in\{0,1\}^{L\times N}$, a list of the $L$ most likely trajectories in the state space (line 7, Fig. 4.(b), and Algorithm 5) according to MAP decision based on the given prior information. Each trajectory is a sequence of $N$ states, representing items classified as either defective or non-defective. Thus, $\mathbf{Z}$ provides $L$ estimations of $\mathbf{U}$. In practice, the $L$ estimations may include any number of defective items and require further processing. The list size parameter $L$ controls the tradeoff between recovery performance and computational complexity. Larger values of $L$ increase the probability that the true defective pattern is contained in the list of candidate trajectories, at the cost of increased computation in subsequent candidate enumeration and MAP selection. This tunable tradeoff is a key advantage of using LVA in the proposed decoding architecture.
@@
In the third step, we extract candidates for the defective set out of the $L$ estimated sequences $\mathbf{Z}$ (lines 9 to 14). For some $l\in\{1,\ldots,L\}$, let $\mathbf{V}^{(l)}$ denote the set of items estimated as defective in $\mathbf{Z}_{l}$, the $l$-th row of $\mathbf{Z}$ (line 11). We ignore sequences that contain fewer than $K_{\min}$ defective items or more than ${\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}$ defective items, for some $\gamma\geq 1$, and consider only $\mathbf{Z}_{l}$ in which ${\color[rgb]{0,0,0}K_{\min}\leq}\left|\mathbf{V}^{(l)}\right|<{\color[rgb]{0,0,0}\lfloor}\gamma{\color[rgb]{0,0,0}K_{\max}{\color[rgb]{0,0,0}\rfloor}}$ as valid sequences. For each one of the valid sequences, we refer to all the combinations of size $k$ in $\mathbf{V}^{(l)}$, for every $k\in\{K_{\min},\ldots,\min(K_{\max},|\mathbf{V}^{(l)}|)\}$, as candidates for $\hat{\mathcal{K}}$ and add them to the candidates list $\mathbf{C}$ (line 12). In particular, for a given set of indices $\mathbf{V}^{(l)}\subseteq\mathcal{N}$ and integers $1\leq K_{\min}\leq K_{\max}$, we define $\mathbf{getAllCombinations}(\mathbf{V}^{(l)},K_{\min},K_{\max})\;\triangleq\;\bigcup_{k=K_{\min}}^{\min(K_{\max},|\mathbf{V}^{(l)}|)}\left\{\mathcal{S}\subseteq\mathbf{V}^{(l)}\;:\;|\mathcal{S}|=k\right\}$. Thus, $\mathbf{getAllCombinations}$ $(\mathbf{V}^{(l)},K_{\min},K_{\max})$ returns the collection of all subsets of $\mathbf{V}^{(l)}$ whose cardinality lies in the range $K_{\min}$ and $K_{\max}$. The operator $\mathbf{getAllCombinations}(\cdot)$ as detailed in Algorithm 6 is used here to generate candidate defective sets without assuming knowledge of the exact sparsity level.
@@
At this point, we have in $\mathbf{C}$ a list of sets of size $k\in\{K_{\min},\ldots,K_{\max}\}$, that are candidates to be our final estimation $\hat{\mathcal{K}}$, and we can calculate the probability of each one of them using $\color[rgb]{0,0,0}\mathbf{\pi},\{\mathbf{\Phi}_{i}\}_{i=\tau+1}^{N}$} (see Algorithm 4). Then, in the fourth step, the estimated defective set, $\hat{\mathcal{K}}$, is finally chosen using MAP estimator out of the $\mathbf{C}$ (line 15), i.e., $\hat{\mathcal{K}}=\arg\max_{c\in\mathbf{C}}{P\left(Y\lvert\mathbf{X},c\right)P\left(c\right)}$ (see Algorithm 7).
@@
If there are no valid sequences in $\mathbf{Z}$, we consider trajectories with fewer than $K_{\min}$ detections for partial recovery. We select the trajectory with the most detections and randomly complete it to form a set of size $K_{\min}$ for our final estimation $\hat{\mathcal{K}}$.
@@
Figure 4: Stage 2 of 2SDGT. (a) All the possible transitions in the state space that we consider in the LVA step, following the insights obtained in Stage 1. These transitions aggregate to a total of 6 trajectories. (b) The two most likely trajectories returned by LVA (assuming $L=2$). Given $K=2$, the black trajectory corresponds to a valid population vector $\mathbf{U}$ with 2 defective items, while the gray trajectory indicates an invalid population with 3 defective items instead. Consequently, in the subsequent step, 2SDGT will extract two optional defective sets: $\left\{U_{6},U_{8}\right\}$ and $\left\{U_{7},U_{8}\right\}$, and will finally choose the most likely one using MAP estimator. (c) Comparison of Stage 2 to ML. With $T=5$, we use the first 5 rows of the testing matrix, ignoring the last test result. This leaves 3 possibly defective items, forming two potentially defective sets of size $K=2$. Using ML, one set is chosen randomly, leading to an error probability of 0.5. With $T=6$, based on the third and sixth test results, there is only one set of size $K=2$ that matches the outcome $Y$, resulting in successful decoding with the ML decoder. As shown above, 2SDGT’s Stage 2 can successfully decode $\mathbf{U}$ with just $T=5$, as using the LVA step it narrows down to only 2 possible trajectories, and then the final estimation is selected based on the given prior information and the insights gained in Stage 1.
@@ III-B Analytical Results
In this section, we provide analytical results related to the proposed 2SDGT. To support our theoretical results, we demonstrate the derived bounds and compare them to simulation results in Fig. 5.
In the correlated-prior model, the number of defectives $K=\sum_{i=1}^{N}U_{i}$ is random but is assumed to satisfy the known bound $K_{\min}\leq K\leq K_{\max}$ (see Section II). Accordingly, throughout this section we analyze Stage 1 under a Bernoulli design with parameter $p=\ln(2)/K_{\max}$ and $T$ tests, where $K_{\max}$ is the design upper bound. Theorem 1 bounds the expected number of items whose status remains unknown after the DND step (equivalently, it bounds the expected number of non-defectives that survive as “possibly defective”), as a function of the realized $K=k$, and we denote the corresponding average contribution by $P_{e,a}^{(DND)}$. Complementarily, Theorem 2 upper bounds the probability that this number exceeds its expectation by a margin, and we denote this upper-tail deviation term by $P_{e,u}^{(DND)}$. These bounds control the size of the remaining search space passed to Stage 2, and in particular ensure that the Stage 2 list-size budget (parameterized by $\gamma$ and the design upper bound $K_{\max}$) is met.
Theorem 1 ([60]).
Consider a group test with a Bernoulli testing matrix with $p=\ln{2}/K_{\max}$ and $T$ tests, and suppose the (random) number of defectives $K$ satisfies $K_{\min}\leq K\leq K_{\max}$. Fix any realization $K=k$ and let $P_{e,a}^{(DND)}\triangleq N^{-\alpha\left(1-\ln{2}/K_{\max}\right)/2}$ for $\alpha\triangleq T/(K_{\max}\log_{2}N)$. The expected number of possibly defective items is bounded by
$\mathbb{E}\left[\left|\mathcal{P}^{(S_{1,1})}\right|\,\big|\,K={\color[rgb]{0,0,0}k}\right]\leq{\color[rgb]{0,0,0}k+\left(N-k\right)}P_{e,a}^{(DND)}.$
The proof of Theorem 1 follows directly from the definition of the DND decoding rule and the Bernoulli test design [22, 60]. A non-defective item is eliminated unless it appears only in tests whose outcomes are positive; equivalently, it survives the DND step if and only if all tests in which it participates are occluded by at least one defective item. Since each item participates in tests independently with probability $p$, the survival event can be characterized by conditioning on the number of tests involving the item and on whether those tests include at least one defective. In particular, under a Bernoulli(p)(p) design with $k$ defectives, $(1-p)^{k}$ denotes the probability that a test is negative, and $p(1-p)^{k-1}$ denotes the probability that a test contains defective $i$ and no other defectives. This leads to the stated expression for the probability of a non-defective item surviving the DND step, which is then used in the subsequent expectation and concentration analysis.
@@
According to [10, Lemma 2], conditioning on $\{K={\color[rgb]{0,0,0}k}\}$ yields
$\mathbb{E}\!\left[\left|\mathcal{P}^{(S_{1,1})}\right|\,\big|\,K={\color[rgb]{0,0,0}k}\right]={\color[rgb]{0,0,0}k+\left(N-k\right)}P_{e}^{(\mathrm{DND})},\quad\text{where}\quad P_{e}^{(\mathrm{DND})}=\bigl(1-p(1-p)^{{\color[rgb]{0,0,0}k}}\bigr)^{T}.$ (2)
By substituting the design choice $p=\ln 2/K_{\max}$, we bound $P_{e}^{(\mathrm{DND})}$ as follows
$P_{e}^{(\mathrm{DND})}$ $=\left(1-p(1-p)^{{\color[rgb]{0,0,0}k}}\right)^{T}$
$=\left(1-p(1-p)^{{\color[rgb]{0,0,0}k-1}}(1-p)\right)^{T}$
$\leq\left(1-p\,e^{-\ln 2}\,(1-p)\right)^{T},$
where the last step follows since $e^{-\ln 2}\leq(1-\ln 2/m)^{m-1}$ for any integer $m\geq 0$, such that, together with the bound ${\color[rgb]{0,0,0}k}\leq K_{\max}$ and with $p=\ln 2/K_{\max}$, we have
$(1-p)^{{\color[rgb]{0,0,0}k-1}}\;\geq\;(1-p)^{K_{\max}-1}\;\geq\;e^{-\ln 2}\;=\;\frac{1}{2},$
as the first inequality holds since 0<1p<10<1-p<1 and ${\color[rgb]{0,0,0}k-1}\leq K_{\max}-1$, and the second from the exponential bound.
Now, applying $(1-z)^{T}\leq e^{-zT}$ with
$z\triangleq p\cdot\frac{1}{2}\cdot(1-p)=\frac{\ln 2}{K_{\max}}\cdot\frac{1}{2}\cdot\left(1-\frac{\ln 2}{K_{\max}}\right),\quad\text{and}\quad T=\alpha K_{\max}\log_{2}N,$
we obtain
$P_{e}^{(\mathrm{DND})}\leq\exp\!\left(-z\cdot\alpha K_{\max}\log_{2}N\right)=N^{-\frac{\alpha}{2}\left(1-\frac{\ln 2}{K_{\max}}\right)}.$
Finally, substituting $P_{e,a}^{(\mathrm{DND})}\triangleq N^{-\frac{\alpha}{2}\left(1-\frac{\ln 2}{K_{\max}}\right)}$ into (2) yields
$\mathbb{E}\!\left[\left|\mathcal{P}^{(S_{1,1})}\right|\,\big|\,K={\color[rgb]{0,0,0}k}\right]\leq{\color[rgb]{0,0,0}k+(N-k)}\,P_{e,a}^{(\mathrm{DND})}.$
@@
Consider a group test with $T$ tests, a Bernoulli testing matrix with the design parameter $p=\ln(2)/K_{\max}$, and the DND decoder. Fix any realization $K=k\in[K_{\min},K_{\max}]$ and let $G$ denote the number of non-defective items marked as possibly defective by DND. Then, for any threshold $g>0$,
$P_{e,u}^{(\mathrm{DND})}(g)\;\triangleq\;P\!\left(G-\mathbb{E}[G]>g\,\big|\,K={\color[rgb]{0,0,0}k}\right)\;\leq\;\frac{\mathrm{Var}(G\mid K={\color[rgb]{0,0,0}k})}{\mathrm{Var}(G\mid K={\color[rgb]{0,0,0}k})+g^{2}},$ (3)
where
$\mathrm{Var}(G\mid K={\color[rgb]{0,0,0}k})=(N-{\color[rgb]{0,0,0}k})(\bar{r}_{1}-\bar{r}_{2})\;+\;(N-{\color[rgb]{0,0,0}k})^{2}(\bar{r}_{2}-\bar{r}_{1}^{\,2}),$ (4)
with $\bar{r}_{1}=\big(1-pq_{0}\big)^{T}$ and $\bar{r}_{2}=\big(1-(2p-p^{2})q_{0}\big)^{T}$ for $q_{0}\triangleq(1-p)^{{\color[rgb]{0,0,0}k}}$.